Construct a higher order models until you are satisfied you have found the best polynomial model.

  1. Write out your final model or display the analysis window for it so I can see the model you settled on.
  2. What is R2-adjusted?
  3. Include the Plot of Fitted Model and Residuals versus X output.
  4. Predict the highway mileage of a car with 150 horsepower. Also construct 95% confidence and prediction intervals for such a car.
  5. Horsepower   MPG Highway
    140   31
    200   25
    172   26
    172   26
    208   30
    110   31
    170   28
    180   25
    170   27
    200   25
    295   25
    110   36
    110   34
    160   28
    110   29
    170   23
    165   20
    170   26
    300   25
    153   28
    141   28
    147   26
    92   33
    93   29
    100   27
    142   21
    100   27
    300   24
    92   33
    214   28
    63   33
    127   30
    96   27
    105   29
    115   30
    145   20
    140   30
    190   26
    55   50
    90   36
    160   31
    102   46
    140   31
    81   33
    124   29
    92   34
    128   27
    278   22
    185   24
    225   23
    160   26
    210   26
    82   37
    103   36
    164   34
    155   24
    255   25
    130   29
    217   25
    100   26
    140   26
    92   33
    202   24
    110   33
    150   30
    151   23
    160   26
    155   31
    110   31
    170   23
    170   28
    92   30
    74   41
    110   31
    160   28
    200   27
    170   28
    140   26
    85   38
    73   37
    90   30
    130   30
    70   43
    82   37
    135   32
    130   29
    138   22
    81   33
    109   21
    134   30
    178   25
    114   28
    168   28
Answer #1

a)

Regression Analysis: MPG Highway versus Horsepower The regression equation is MPG Highway = 38.1 -0.0630 Horsepower Predictor

From the above three polynomial models with degrees 1, 2 and 3 respectively on the explanatory variable  Horsepower, the corresponding R-sq(adj) are 37.6%, 50.1%, and 53.7% respectively. The increase in R-sq(adj) from degree 1 to degree 2 model is approximately 10 with an increasing 1 parameter on the model. Hence, the model 2 with degree 2 fits better than model 1 with degree 1. The increase in R-sq(adj) from degree 2 to degree 3 model is approximately 4 with an increasing 1 parameter on the model. Hence, the model 3 with degree 3 does not increase the fitness fits than model 2. Hence, the best polynomial model for these given data is the model with degree 2.

b) The R2-adjusted= R-Sq(adj) = 50.1%

c)  the Plot of Fitted Model and Residuals versus X output is

Scatterplot of Residuals vs Fitted Residuals 25.0 27.5 30.0 32.5 Fitted 35.0 37.5 40.0

From the above residual VS fitted value plot, the plot shows that the residuals and the fitted values are uncorrelated. Also, the linear model should be in a homoscedastic with normally distributed errors. Whereas, some of the points (around three points) have deviated largely from the remaining. Hence, some data points may be the outlier.

d) Predict the highway mileage of a car with 150 horsepower. Also, construct 95% confidence and prediction intervals for such a car.

Ans: The predicted highway mileage of a car with 150 horsepower is

MPG Highway = 50.2 - 0.225* 150 + 0.000482 *150^2=  27.262.

95% confidence intervals for such a car is  (26.286, 28.239).

95% prediction intervals for such a car is  (19.715, 34.809).